OpenStudy (anonymous):

What are the real values of x such that f(x)=0 f(x)=9xsquare - 25

OpenStudy (anonymous):

\(\Huge 9x^2-25=0 \) that can be written as the difference of squares... which can be factored.... can you factor that left hand side?

OpenStudy (anonymous):

I can't.

OpenStudy (anonymous):

\(\Huge 9x^2-25=0 \) \(\Huge (3x)^2-5^2=0 \) now try to factor using difference of squares.....

OpenStudy (anonymous):

use this cheat sheet.... very handy.... it's the first one on page 2 in "factoring".... http://tutorial.math.lamar.edu/pdf/Algebra_Cheat_Sheet.pdf

OpenStudy (anonymous):

The real values are x= -5/3 and 5/3??